OpenStudy (anonymous):

how many real solutions are there ? 4x2 + 16x + 16 = 0

OpenStudy (idealist10):

4x^2+16x+16=0 factor out the 4 4(x^2+4x+4)=0 x^2+4x+4=0 factor (x+2)(x+2)=0 x+2=0 x=-2 Answer: One solution